how much 60% solution must be mixed with a 80% solution to produce 25 gallons of a 76% solution
Answer by mananth(16946)

how much 60% solution let it be x gallons
80% solution 25-x
mix gallons of a 76% solution--- 25
..
0.6x+0.8(25-x)=0.76*25
0.6x+20-0.8x=19
-0.2x=-1
x=-1/-0.2
x= 5 gallons of 60% solution has to be added.
...
m.ananth@hotmail.ca